myininaya (myininaya):

find the three consecutive integers such that the square of the largest exceeds the sum of the quarters of the squares of the other two by 12

myininaya (myininaya):

i don't get it i cannot get an integer x x+2 x+4 (x+4)^2=1/4*x^2+1/4*(x+2)^2+12 is what i did and got x^2+14x+6=0

myininaya (myininaya):

i cannot get an integer from x^2+14x+6=0

OpenStudy (radar):

Let x = first integer then x+1 = second integer and x+2 = third integer ((x+1)/4)^2 +(x^2)/4=(x+2)^2 + 12

OpenStudy (radar):

Problem does not say even or odd just consecutive

myininaya (myininaya):

oops i mean even intgers sorry

myininaya (myininaya):

find the first three consecutive even integers

OpenStudy (radar):

O.K that does change things lol

myininaya (myininaya):

ignore the first three lol i was trying to do this problem for someone else who asked it earlier but the above is the only way i could translate it

OpenStudy (radar):

Let 2x = first even integer then 2x+2 = second and 2x+4 = third

OpenStudy (anonymous):

no such integers exist

myininaya (myininaya):

thats what i was thinking so why would they asked this question

OpenStudy (anonymous):

either you typed the question out incorrectly or they didn't check their solutions thoroughly

myininaya (myininaya):

well there is no way for me to know if the question is typed incorrectly or not i was trying to answer this same question earlier from someone else who posted it on openstudy i typed it exactly as he did

myininaya (myininaya):

without the word even lol

OpenStudy (anonymous):

ah, okay. it will suffice to say there's no solution

myininaya (myininaya):

i just thought they wouldn't try to trick pre algebra peeps so i was like i had to be misunderstanding or the question was typed incorrectly thanks jamesm

OpenStudy (radar):

My solution was not an integer. I can't work it!

myininaya (myininaya):

thanks radar

OpenStudy (radar):

I too, wonder why such a problem was assigned, sure doesn't reinforce confidence!

myininaya (myininaya):

i think it is possible that the person did not write it correctly

OpenStudy (radar):

Ah so.

myininaya (myininaya):

but i did ask if he wrote it word from word and he said yes so maybe not idk

OpenStudy (anonymous):

is it possible that they meant 15 and not 12? 15 yields integer solutions

myininaya (myininaya):

hmm.. maybe he is still here i will ask about the 15